OpenStudy (asad786):

@StudyGurl14 need help

OpenStudy (asad786):

@StudyGurl14

OpenStudy (studygurl14):

Okay, so you want to know how to find the area of a circle, correct?

OpenStudy (asad786):

yes

OpenStudy (asad786):

hello

OpenStudy (asad786):

take that same question how did he do it exactly

OpenStudy (studygurl14):

Okay, so the formula of the area of a circle is \(\large r^2\pi\) r = radius

OpenStudy (asad786):

whats the 2 for

OpenStudy (studygurl14):

That question said to find the area of a round glass table that has a diameter of 24

OpenStudy (asad786):

yes

OpenStudy (studygurl14):

Diameter is 2 times the radius, so to find the radius, you divide the diameter by 2. 24/2 = 12 radius = 12

OpenStudy (asad786):

yes

OpenStudy (studygurl14):

So, using that information, and the formula, can you find the area?

OpenStudy (asad786):

12 times 3.14

OpenStudy (asad786):

right

OpenStudy (studygurl14):

No. You keep forgetting the square part.

OpenStudy (studygurl14):

It is the radius SQUARED

OpenStudy (studygurl14):

How old are you?

OpenStudy (asad786):

12

OpenStudy (asad786):

whats the square

OpenStudy (studygurl14):

Have you ever heard of exponents?

OpenStudy (asad786):

yes

OpenStudy (asad786):

its in pemdas

OpenStudy (studygurl14):

That's what the 2 is. It's an exponent of 2

OpenStudy (asad786):

oh 12 times 12 times 3.14

OpenStudy (studygurl14):

RIGHT

OpenStudy (asad786):

oh thanks for the help

OpenStudy (studygurl14):

You're welcome. :)